OpenStudy (anonymous):

A model scale is 1 cm = 8 m. If the actual object is 40 meters, the model is______ a0 cm long. help me please:(

OpenStudy (anonymous):

For every 1cm there is 8m. So how many 8m is in 40m?

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

8m -------------- 1cm 1m ----------------- \(\large \frac{1}{8}\) cm 40m -----------------\(\large \frac{1}{8} \times 40 = ??\) cm

OpenStudy (anonymous):

is it 80?

OpenStudy (anonymous):

How?? \[\frac{1}{8} \times 40 = How \; \; Much??\]

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

heelp